# Implicit Regularization of Infinitesimally-perturbed Gradient Descent Toward Low-dimensional Solutions

### v1

@@ 6 Application: Over-parameterized Matrix Sensing
To proceed, we decompose ${\mathbf{M}^{\star}}=\mathbf{V}^{\star}\bm{\Sigma}^{\star}\mathbf{V}^{\star\top}$ as
${\mathbf{M}^{\star}}=\underbrace{\mathbf{V}_{\mathbf{X}}\bm{\Sigma}^{\star}\mathbf{V}^{\top}_{\mathbf{X}}}_{:=\mathbf{M}^{\star}_{\mathbf{X}}}+\underbrace{\left(\mathbf{V}^{\star}-\mathbf{V}_{\mathbf{X}}\right)\bm{\Sigma}^{\star}\mathbf{V}^{\star\top}+\mathbf{V}_{\mathbf{X}}\bm{\Sigma}^{\star}\left(\mathbf{V}^{\star}-\mathbf{V}_{\mathbf{X}}\right)^{\top}}_{:=\mathbf{\Delta_{M}}},$ (95)
[…]
$\quad+\underbrace{\left\lVert\mathcal{P}_{\mathbf{X}}\left(\mathbf{I}-\eta\mathbf{X}\mathbf{X}^{\top}\right)\left(\left(\mathbf{I}+\eta\left(\mathbf{M}^{\star}_{\mathbf{X}}-\mathbf{X}\mathbf{X}^{\top}\right)\right)^{-1}-\left(\mathbf{I}+\eta\left({\mathbf{M}^{\star}}-\mathbf{X}\mathbf{X}^{\top}\right)\right)^{-1}\right)\right\rVert}_{(\textup{III}_{2})}.$
@@
$(\textup{III}_{1})$ $=\left\lVert\left(\mathbf{I}-\eta\bm{\Sigma}_{\mathbf{X}}^{2}\right)\left(\mathbf{I}+\eta\left(\bm{\Sigma}^{\star}-\bm{\Sigma}_{\mathbf{X}}^{2}\right)\right)^{-1}\right\rVert$ (97)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left\lVert\left(\mathbf{I}+\eta\bm{\Sigma}^{\star}\right)^{-1}\right\rVert$
$=\frac{1}{1+\eta\sigma_{r}^{\star}}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}1-\frac{1}{2}\eta\sigma_{r}^{\star}.$
@@
To control $(\textup{III}_{2})$, we rely on Lemma 27. Upon setting $\mathbf{A}=\mathbf{I}+\eta\left(\mathbf{M}^{\star}_{\mathbf{X}}-\mathbf{X}\mathbf{X}^{\top}\right)$ and $\mathbf{B}=\mathbf{I}+\eta\left({\mathbf{M}^{\star}}-\mathbf{X}\mathbf{X}^{\top}\right)$ in Lemma 27, we have
@@
$\left\lVert\mathbf{A}-\mathbf{B}\right\rVert$ $=\eta\left\lVert\mathbf{\Delta_{M}}\right\rVert\leq 2\sigma_{1}^{\star}\left\lVert\mathbf{V}^{\star}-\mathbf{V}_{\mathbf{X}}\right\rVert\leq 2\sqrt{2}\sigma_{1}^{\star}\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert.$ (100)
@@
[…]
$=\eta\left\lVert(\mathcal{A}^{*}\mathcal{A}-I)(\bm{\Delta})\left(\mathbf{I}+\eta\bm{\Delta}\right)^{-1}\tilde{\mathbf{X}}^{+}\big{(}\tilde{\mathbf{X}}^{+}\big{)}^{\dagger}\right\rVert$
$\leq\eta\left\lVert(\mathcal{A}^{*}\mathcal{A}-I)\left(\bm{\Delta}\right)\right\rVert\left\lVert\left(\mathbf{I}+\eta\bm{\Delta}\right)^{-1}\right\rVert$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\eta\sqrt{2r}\delta\left\lVert\bm{\Delta}\right\rVert\left\lVert\left(\mathbf{I}+\eta\bm{\Delta}\right)^{-1}\right\rVert$
$\stackrel{{\scriptstyle(c)}}{{\leq}}4\eta\sqrt{r}\sigma_{1}^{\star}\delta,$
[…]
@@
where the last inequality follows from $\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert\leq 4\sqrt{r}\kappa\delta$ by our assumption. This completes the proof. ∎
$\mathcal{M}$-strict saddle property and $\mathcal{M}$-regularity property
Next, we establish the $\mathcal{M}$-strict saddle property and the $\mathcal{M}$-regularity property, both of which are essential for ensuring the efficient convergence of our algorithm.
@@
  • It satisfies $\left(\frac{1}{160}\sigma_{r}^{\star 3/2},\frac{1}{5}\sigma_{r}^{\star},\frac{1}{8}\sigma_{r}^{\star 1/2}\right)$$-\mathcal{M}$-strict saddle property (Definition 6).
  • It satisfies $\left(\frac{1}{2}\sigma_{r}^{\star},\frac{425}{64}\sigma_{1}^{\star},\frac{1}{4}\sigma_{r}^{\star 1/2}\right)$$-\mathcal{M}$-regularity property (Definition 7).
To prove the above proposition, we establish a connection between the local geometry of $f(\mathbf{X})$ within the implicit region $\mathcal{M}$ to that of the exactly-parameterized matrix sensing problem, defined as
@@
provided that the linear operator $\mathcal{A}$ satisfies $(\delta,6r)$-RIP with $\delta\leq\frac{1}{10}$ [42, Theorem 8]. [^6] Note that the definition of ${\mathcal{X}}^{\star}_{\leq r}$ should not be confused with the previously introduced set $\mathcal{X}^{\star}=\{\mathbf{X}:\mathbf{X}=\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}^{\top},\mathbf{O}\in\mathcal{O}_{r^{\prime}\times r}\}$, which characterizes the true solutions in the over-parameterized space. Moreover, this objective function satisfies the strict saddle property and exhibits strong local structure globally (i.e., over the exactly-parameterized space $\mathbb{R}^{n\times r}$), as detailed below.
@@
  • It satisfies $\left(\epsilon,\frac{1}{5}\sigma_{r}^{\star},\frac{20\epsilon}{\sigma_{r}^{\star}}\right)$-strict saddle property, for any $\epsilon>0$ [42, Theorem 8].
  • It satisfies $\left(\frac{1}{2}\sigma_{r}^{\star},\frac{425}{64}\sigma_{1}^{\star},\frac{1}{4}\sigma_{r}^{\star 1/2}\right)$-regularity property [59, Eq. 5.7].
To prove Proposition 28, we leverage the above lemma to show that, when restricted to the implicit region $\mathcal{M}$ (as defined in Equation 78), the function $f(\mathbf{X})$ satisfies the same properties with exactly the same parameters.
@@
We first establish a one-to-one mapping $h:\mathcal{M}\to\mathbb{R}^{n\times r}$, which maps an arbitrary $\mathbf{X}\in\mathcal{M}$ to a point ¯$\bar{}\mathbf{X}=h(\mathbf{X})\in\mathbb{R}^{n\times r}$. Specifically, for an arbitrary $\mathbf{X}\in\mathcal{M}$ with the SVD of the form $\mathbf{X}=\mathbf{L}_{\mathbf{X}}\bm{\Sigma}_{\mathbf{X}}\mathbf{R}_{\mathbf{X}}^{\top}$, where $\mathbf{L}_{\mathbf{X}}\in\mathcal{O}_{n\times r}$, $\mathbf{R}_{\mathbf{X}}\in\mathcal{O}_{r^{\prime}\times r}$, and $\bm{\Sigma}_{\mathbf{X}}\in\mathbb{R}^{r\times r}$ is a diagonal matrix [^7] If $\mathrm{rank}(\mathbf{X})<r$, we construct $\bm{\Sigma}_{\mathbf{X}}$ by padding its diagonal with additional zero entries., we define the mapping as $\bar{\mathbf{X}}=h(\mathbf{X})=\mathbf{L}_{\mathbf{X}}\bm{\Sigma}_{\mathbf{X}}\in\mathbb{R}^{n\times r}$.
@@
Suppose that $\mathbf{X}\in\mathcal{M}$ satisfies $\left\lVert\nabla f(\mathbf{X})\right\rVert_{\mathrm{F}}<\epsilon$ and $\operatorname{dist}(\mathbf{X},\mathcal{X}_{\mathcal{M}})>\frac{20\epsilon}{\sigma_{r}^{\star}}$. To establish the $\mathcal{M}$-strict saddle property with the specified parameters, it suffices to show that $\lambda_{\min}\left(\nabla^{2}f(\mathbf{X})\right)\leq-\frac{1}{5}\sigma_{r}^{\star}$. To this end, we first show that for $\bar{\mathbf{X}}=h(\mathbf{X})$, we also have $\left\lVert\nabla\bar{f}\left(\bar{\mathbf{X}}\right)\right\rVert_{\mathrm{F}}<\epsilon$ and $\operatorname{dist}\left(\bar{\mathbf{X}},{\mathcal{X}}^{\star}_{\leq r}\right)>\frac{20\epsilon}{\sigma_{r}^{\star}}$. One can write
$\left\lVert\nabla f(\mathbf{X})\right\rVert_{\mathrm{F}}$ $=\left\lVert\mathcal{A}^{*}\mathcal{A}\big{(}\mathbf{X}\mathbf{X}^{\top}-{\mathbf{M}^{\star}}\big{)}\mathbf{X}\right\rVert_{\mathrm{F}}$ (105)
$=\left\lVert\mathcal{A}^{*}\mathcal{A}\left(\bar{\mathbf{X}}\bar{\mathbf{X}}^{\top}-{\mathbf{M}^{\star}}\right)\bar{\mathbf{X}}\mathbf{R}_{\mathbf{X}}^{\top}\right\rVert_{\mathrm{F}}$
$=\left\lVert\mathcal{A}^{*}\mathcal{A}\left(\bar{\mathbf{X}}\bar{\mathbf{X}}^{\top}-{\mathbf{M}^{\star}}\right)\bar{\mathbf{X}}\right\rVert_{\mathrm{F}}$
$=\left\lVert\nabla\bar{f}\left(\bar{\mathbf{X}}\right)\right\rVert_{\mathrm{F}},$
@@
$\operatorname{dist}(\mathbf{X},\mathcal{X}_{\mathcal{M}})$ $=\min_{\mathbf{O}\in\mathcal{O}_{r^{\prime}\times r^{\prime}}}\left\lVert\mathbf{X}\mathbf{O}-\begin{bmatrix}\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}&\bm{0}_{n\times(r^{\prime}-r)}\end{bmatrix}\right\rVert_{\mathrm{F}}$ (106)
@@
$=\min_{\bar{}\mathbf{O}\in\mathcal{O}_{r\times r}}\left\lVert\bar{\mathbf{X}}\bar{}\mathbf{O}-\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\right\rVert_{\mathrm{F}}$
$=\operatorname{dist}\left(\bar{\mathbf{X}},{\mathcal{X}}^{\star}_{\leq r}\right),$
where $(a)$ follows after setting $\mathbf{O}=\begin{bmatrix}\mathbf{R}_{\mathbf{X}}\bar{}\mathbf{O}&\mathbf{R}_{\mathbf{X}}^{\perp}\end{bmatrix}\in\mathcal{O}_{r^{\prime}\times r^{\prime}}$ and choosing $\mathbf{R}_{\mathbf{X}}^{\perp}\in\mathcal{O}_{n\times(r^{\prime}-r)}$ that satisfies $\mathbf{R}_{\mathbf{X}}^{\top}\mathbf{R}_{\mathbf{X}}^{\perp}=\bm{0}$. This implies that $\operatorname{dist}\left(\bar{\mathbf{X}},{\mathcal{X}}^{\star}_{\leq r}\right)=\operatorname{dist}(\mathbf{X},\mathcal{X}_{\mathcal{M}})>\frac{20\epsilon}{\sigma_{r}^{\star}}$. Therefore, we have established that $\left\lVert\nabla f(\mathbf{X})\right\rVert_{\mathrm{F}}\leq\epsilon$ and $\operatorname{dist}\left(\bar{\mathbf{X}},{\mathcal{X}}^{\star}_{\leq r}\right)>\frac{20\epsilon}{\sigma_{r}^{\star}}$. Based on these inequalities, the first statement of Lemma 29 implies that $\lambda_{\min}\left(\nabla^{2}\bar{f}\left(\bar{\mathbf{X}}\right)\right)\leq-\frac{1}{5}\sigma_{r}^{\star}$. Next, we use this inequality to establish a similar bound for $\lambda_{\min}\left(\nabla^{2}f(\mathbf{X})\right)$. This is achieved as follows:
@@
$=\lambda_{\min}\left(\nabla^{2}\bar{f}\left(\bar{\mathbf{X}}\right)\right)$
$\leq-\frac{1}{5}\sigma_{r}^{\star},$
where $(a)$ follows after setting $\mathbf{Z}=\bar{\mathbf{Z}}\mathbf{R}_{\mathbf{X}}^{\top}$. Therefore, $f(\mathbf{X})$ satisfies $\left(\epsilon,\frac{1}{5}\sigma_{r}^{\star},\frac{20\epsilon}{\sigma_{r}^{\star}}\right)$$-\mathcal{M}$-strict saddle property. Setting $\epsilon=\frac{1}{160}\sigma_{r}^{\star 3/2}$ completes the proof of the first statement of Proposition 28.

### v2

@@ 6 Application: Over-parameterized Matrix Sensing
$\stackrel{{\scriptstyle(a)}}{{=}}\mathcal{P}_{\mathbf{V}^{\star\perp}}\left(\mathbf{I}-\eta\mathbf{X}\mathbf{X}^{\top}\right)\mathbf{X}(\tilde{\mathbf{X}}^{+})^{\dagger}\stackrel{{\scriptstyle(b)}}{{=}}\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\left(\mathbf{I}-\eta\mathbf{X}\mathbf{X}^{\top}\right)(\mathbf{I}+\eta\bm{\Delta})^{-1}\mathcal{P}_{\tilde{\mathbf{X}}^{+}}$,
where, in $(a)$, we use the fact that $\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathbf{\Theta}^{\star}=\bm{0}$. Moreover, in $(b)$, we use the fact that $\mathbf{X}(\tilde{\mathbf{X}}^{+})^{\dagger}=(\mathbf{I}+\eta\bm{\Delta})^{-1}\tilde{\mathbf{X}}^{+}(\tilde{\mathbf{X}}^{+})^{\dagger}=(\mathbf{I}+\eta\bm{\Delta})^{-1}\mathcal{P}_{\tilde{\mathbf{X}}^{+}}$. Therefore, we obtain
$(\textup{I})$ $\leq\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert\underbrace{\left\lVert\mathcal{P}_{\mathbf{X}}\left(\mathbf{I}-\eta\mathbf{X}\mathbf{X}^{\top}\right)\left(\mathbf{I}+\eta\bm{\Delta}\right)^{-1}\right\rVert}_{(\textup{III})}.$ (95)
To proceed, we decompose $\mathbf{\Theta}^{\star}=\mathbf{V}^{\star}\bm{\Sigma}^{\star}\mathbf{V}^{\star\top}$ as
$\mathbf{\Theta}^{\star}=\underbrace{\mathbf{V}_{\mathbf{X}}\bm{\Sigma}^{\star}\mathbf{V}^{\top}_{\mathbf{X}}}_{:=\mathbf{\Theta}^{\star}_{\mathbf{X}}}+\underbrace{\left(\mathbf{V}^{\star}-\mathbf{V}_{\mathbf{X}}\right)\bm{\Sigma}^{\star}\mathbf{V}^{\star\top}+\mathbf{V}_{\mathbf{X}}\bm{\Sigma}^{\star}\left(\mathbf{V}^{\star}-\mathbf{V}_{\mathbf{X}}\right)^{\top}}_{:=\mathbf{\Delta_{M}}},$ (96)
where $\mathbf{V}_{\mathbf{X}}$ is defined as $\mathbf{V}_{\mathbf{X}}=\operatorname*{arg\,min}_{\mathbf{V}\in\mathcal{O}_{n\times r},\mathcal{P}_{\mathbf{X}}\mathbf{V}=\mathbf{V}}\left\lVert\mathbf{V}^{\star}-\mathbf{V}\right\rVert$. Therefore, we can decompose $(\textup{III})$ as
$(\textup{III})$ $\leq\underbrace{\left\lVert\mathcal{P}_{\mathbf{X}}\left(\mathbf{I}-\eta\mathbf{X}\mathbf{X}^{\top}\right)\left(\mathbf{I}+\eta\left(\mathbf{\Theta}^{\star}_{\mathbf{X}}-\mathbf{X}\mathbf{X}^{\top}\right)\right)^{-1}\right\rVert}_{(\textup{III}_{1})}$ (97)
$\quad+\underbrace{\left\lVert\mathcal{P}_{\mathbf{X}}\left(\mathbf{I}-\eta\mathbf{X}\mathbf{X}^{\top}\right)\left(\left(\mathbf{I}+\eta\left(\mathbf{\Theta}^{\star}_{\mathbf{X}}-\mathbf{X}\mathbf{X}^{\top}\right)\right)^{-1}-\left(\mathbf{I}+\eta\left(\mathbf{\Theta}^{\star}-\mathbf{X}\mathbf{X}^{\top}\right)\right)^{-1}\right)\right\rVert}_{(\textup{III}_{2})}.$
@@
$(\textup{III}_{1})$ $=\left\lVert\left(\mathbf{I}-\eta\bm{\Sigma}_{\mathbf{X}}^{2}\right)\left(\mathbf{I}+\eta\left(\bm{\Sigma}^{\star}-\bm{\Sigma}_{\mathbf{X}}^{2}\right)\right)^{-1}\right\rVert\stackrel{{\scriptstyle(a)}}{{\leq}}\left\lVert\left(\mathbf{I}+\eta\bm{\Sigma}^{\star}\right)^{-1}\right\rVert=\frac{1}{1+\eta\sigma_{r}^{\star}}\stackrel{{\scriptstyle(b)}}{{\leq}}1-\frac{1}{2}\eta\sigma_{r}^{\star}.$ (98)
@@
To control $(\textup{III}_{2})$, we rely on Lemma 15. Upon setting $\mathbf{A}=\mathbf{I}+\eta\left(\mathbf{\Theta}^{\star}_{\mathbf{X}}-\mathbf{X}\mathbf{X}^{\top}\right)$ and $\mathbf{B}=\mathbf{I}+\eta\left(\mathbf{\Theta}^{\star}-\mathbf{X}\mathbf{X}^{\top}\right)$ in Lemma 15, we have
@@
$\left\lVert\mathbf{A}-\mathbf{B}\right\rVert$ $=\eta\left\lVert\mathbf{\Delta_{M}}\right\rVert\leq 2\sigma_{1}^{\star}\left\lVert\mathbf{V}^{\star}-\mathbf{V}_{\mathbf{X}}\right\rVert\leq 2\sqrt{2}\sigma_{1}^{\star}\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert,$ (101)
@@
$(\textup{I})$ $\leq\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert\left((\textup{III}_{1})+(\textup{III}_{2})\right)\leq\left(1-\frac{1}{2}\eta\sigma_{r}^{\star}+8\sqrt{2}\eta\sigma_{1}^{\star}\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert\right)\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert.$ (102)
Finally, we control $(\textup{II})$. To this goal, we write
@@
$\leq\eta\left\lVert(\mathcal{A}^{*}\mathcal{A}-I)\left(\bm{\Delta}\right)\right\rVert\left\lVert\left(\mathbf{I}+\eta\bm{\Delta}\right)^{-1}\right\rVert\stackrel{{\scriptstyle(b)}}{{\leq}}\eta\sqrt{2r}\delta\left\lVert\bm{\Delta}\right\rVert\left\lVert\left(\mathbf{I}+\eta\bm{\Delta}\right)^{-1}\right\rVert\stackrel{{\scriptstyle(c)}}{{\leq}}4\eta\sqrt{r}\sigma_{1}^{\star}\delta,$
where in $(a)$, we use Lemma 14. Moreover, in $(b)$, we use Lemma 12. Finally, in $(c)$, we use $\left\lVert\bm{\Delta}\right\rVert\leq 4\sigma_{1}^{\star}$ and $\eta\leq\frac{1}{4\sqrt{2}\sigma_{1}^{\star}}$. Combining the derived bounds for $(\textup{I})$ and $(\textup{II})$, we have
@@
where the last inequality follows from $\left\lVert\mathcal{P}_{\mathbf{V}^{\star\perp}}\mathcal{P}_{\mathbf{X}}\right\rVert\leq 4\sqrt{r}\kappa\delta$ by our assumption. ∎
$\mathcal{M}_{\text{RIP}}$-strict saddle property and $\mathcal{M}_{\text{RIP}}$-regularity property
Next, we establish the $\mathcal{M}_{\text{RIP}}$-strict saddle property and the $\mathcal{M}_{\text{RIP}}$-regularity property, both of which are essential for ensuring the efficient convergence of our algorithm.
@@
  • It satisfies $\left(\frac{1}{160}\sigma_{r}^{\star 3/2},\frac{1}{5}\sigma_{r}^{\star},\frac{1}{8}\sigma_{r}^{\star 1/2}\right)$-$\mathcal{M}_{\text{RIP}}$-strict saddle property (Definition 6).
  • It satisfies $\left(\frac{1}{2}\sigma_{r}^{\star},\frac{425}{64}\sigma_{1}^{\star},\frac{1}{4}\sigma_{r}^{\star 1/2}\right)$-$\mathcal{M}_{\text{RIP}}$-regularity property (Definition 7).
To prove the above proposition, we establish a connection between the local geometry of $f(\mathbf{X})$ within the implicit region $\mathcal{M}_{\text{RIP}}$ to that of the exactly-parameterized matrix sensing problem:
@@
provided that the linear operator $\mathcal{A}$ satisfies $(\delta,6r)$-RIP with $\delta\leq\frac{1}{10}$ [ge2017no, Theorem 8]. [^4] Note that the definition of ${\mathcal{X}}^{\star}_{\leq r}$ should not be confused with the previously introduced set $\mathcal{X}^{\star}=\{\mathbf{X}:\mathbf{X}=\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\mathbf{O}^{\top},\mathbf{O}\in\mathcal{O}_{r^{\prime}\times r}\}$, which characterizes the true solutions in the over-parameterized regime. Moreover, this objective function satisfies the strict saddle property and exhibits strong local structure globally (i.e., over the exactly-parameterized space $\mathbb{R}^{n\times r}$), as detailed below.
@@
  • It satisfies $\left(\epsilon,\frac{1}{5}\sigma_{r}^{\star},\frac{20\epsilon}{\sigma_{r}^{\star}}\right)$-strict saddle property, for any $\epsilon>0$ [ge2017no, Theorem 8].
  • It satisfies $\left(\frac{1}{2}\sigma_{r}^{\star},\frac{425}{64}\sigma_{1}^{\star},\frac{1}{4}\sigma_{r}^{\star 1/2}\right)$-regularity property [tu2016low, Eq. 5.7].
To prove Proposition 8, we leverage the above lemma to show that, when restricted to the implicit region $\mathcal{M}_{\text{RIP}}$ (as defined in Equation 79), the function $f(\mathbf{X})$ satisfies the same properties with exactly the same parameters.
@@
We first establish a one-to-one mapping $h:\mathcal{M}_{\text{RIP}}\to\mathbb{R}^{n\times r}$, which maps an arbitrary $\mathbf{X}\in\mathcal{M}_{\text{RIP}}$ to a point $\bar{}\mathbf{X}=h(\mathbf{X})\in\mathbb{R}^{n\times r}$. Specifically, for an arbitrary $\mathbf{X}\in\mathcal{M}_{\text{RIP}}$ with the SVD of the form $\mathbf{X}=\mathbf{L}_{\mathbf{X}}\bm{\Sigma}_{\mathbf{X}}\mathbf{R}_{\mathbf{X}}^{\top}$, where $\mathbf{L}_{\mathbf{X}}\in\mathcal{O}_{n\times r}$, $\mathbf{R}_{\mathbf{X}}\in\mathcal{O}_{r^{\prime}\times r}$, and $\bm{\Sigma}_{\mathbf{X}}\in\mathbb{R}^{r\times r}$ is a diagonal matrix [^5] If $\mathrm{rank}(\mathbf{X})<r$, we construct $\bm{\Sigma}_{\mathbf{X}}$ by padding its diagonal with additional zero entries., we define the mapping as $\bar{\mathbf{X}}=h(\mathbf{X})=\mathbf{L}_{\mathbf{X}}\bm{\Sigma}_{\mathbf{X}}\in\mathbb{R}^{n\times r}$.
@@
Suppose that $\mathbf{X}\in\mathcal{M}_{\text{RIP}}$ satisfies $\left\lVert\nabla f(\mathbf{X})\right\rVert_{\mathrm{F}}<\epsilon$ and $\operatorname{dist}(\mathbf{X},\mathcal{X}_{\mathcal{M}_{\text{RIP}}})>\frac{20\epsilon}{\sigma_{r}^{\star}}$. To establish the $\mathcal{M}_{\text{RIP}}$-strict saddle property with the specified parameters, it suffices to show that $\lambda_{\min}\left(\nabla^{2}f(\mathbf{X})\right)\leq-\frac{1}{5}\sigma_{r}^{\star}$. To this end, we first show that for $\bar{\mathbf{X}}=h(\mathbf{X})$, we also have $\left\lVert\nabla\bar{f}\left(\bar{\mathbf{X}}\right)\right\rVert_{\mathrm{F}}<\epsilon$ and $\operatorname{dist}\left(\bar{\mathbf{X}},{\mathcal{X}}^{\star}_{\leq r}\right)>\frac{20\epsilon}{\sigma_{r}^{\star}}$. One can write
$\left\lVert\nabla f(\mathbf{X})\right\rVert_{\mathrm{F}}$ $=\left\lVert\mathcal{A}^{*}\mathcal{A}\big(\mathbf{X}\mathbf{X}^{\top}-\mathbf{\Theta}^{\star}\big)\mathbf{X}\right\rVert_{\mathrm{F}}=\left\lVert\mathcal{A}^{*}\mathcal{A}\left(\bar{\mathbf{X}}\bar{\mathbf{X}}^{\top}-\mathbf{\Theta}^{\star}\right)\bar{\mathbf{X}}\mathbf{R}_{\mathbf{X}}^{\top}\right\rVert_{\mathrm{F}}$ (106)
$=\left\lVert\mathcal{A}^{*}\mathcal{A}\left(\bar{\mathbf{X}}\bar{\mathbf{X}}^{\top}-\mathbf{\Theta}^{\star}\right)\bar{\mathbf{X}}\right\rVert_{\mathrm{F}}=\left\lVert\nabla\bar{f}\left(\bar{\mathbf{X}}\right)\right\rVert_{\mathrm{F}},$
@@
$\operatorname{dist}(\mathbf{X},\mathcal{X}_{\mathcal{M}_{\text{RIP}}})$ $=\min_{\mathbf{O}\in\mathcal{O}_{r^{\prime}\times r^{\prime}}}\left\lVert\mathbf{X}\mathbf{O}-\begin{bmatrix}\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}&\bm{0}_{n\times(r^{\prime}-r)}\end{bmatrix}\right\rVert_{\mathrm{F}}$ (107)
@@
$=\min_{\bar{}\mathbf{O}\in\mathcal{O}_{r\times r}}\left\lVert\bar{\mathbf{X}}\bar{}\mathbf{O}-\mathbf{V}^{\star}\bm{\Sigma}^{\star 1/2}\right\rVert_{\mathrm{F}}=\operatorname{dist}\left(\bar{\mathbf{X}},{\mathcal{X}}^{\star}_{\leq r}\right),$
where $(a)$ follows after setting $\mathbf{O}=\begin{bmatrix}\mathbf{R}_{\mathbf{X}}\bar{}\mathbf{O}&\mathbf{R}_{\mathbf{X}}^{\perp}\end{bmatrix}\in\mathcal{O}_{r^{\prime}\times r^{\prime}}$ and choosing $\mathbf{R}_{\mathbf{X}}^{\perp}\in\mathcal{O}_{n\times(r^{\prime}-r)}$ that satisfies $\mathbf{R}_{\mathbf{X}}^{\top}\mathbf{R}_{\mathbf{X}}^{\perp}=\bm{0}$. This implies that $\operatorname{dist}\left(\bar{\mathbf{X}},{\mathcal{X}}^{\star}_{\leq r}\right)=\operatorname{dist}(\mathbf{X},\mathcal{X}_{\mathcal{M}_{\text{RIP}}})>\frac{20\epsilon}{\sigma_{r}^{\star}}$. Therefore, we have established that $\left\lVert\nabla f(\mathbf{X})\right\rVert_{\mathrm{F}}\leq\epsilon$ and $\operatorname{dist}\left(\bar{\mathbf{X}},{\mathcal{X}}^{\star}_{\leq r}\right)>\frac{20\epsilon}{\sigma_{r}^{\star}}$. Based on these inequalities, the first statement of Lemma 16 implies that $\lambda_{\min}\left(\nabla^{2}\bar{f}\left(\bar{\mathbf{X}}\right)\right)\leq-\frac{1}{5}\sigma_{r}^{\star}$. Next, we use this inequality to establish a similar bound for $\lambda_{\min}\left(\nabla^{2}f(\mathbf{X})\right)$. This is achieved as follows:
@@
$=\lambda_{\min}\left(\nabla^{2}\bar{f}\left(\bar{\mathbf{X}}\right)\right)\leq-\frac{1}{5}\sigma_{r}^{\star},$
where $(a)$ follows after setting $\mathbf{Z}=\bar{\mathbf{Z}}\mathbf{R}_{\mathbf{X}}^{\top}$. Therefore, $f(\mathbf{X})$ satisfies $\left(\epsilon,\frac{1}{5}\sigma_{r}^{\star},\frac{20\epsilon}{\sigma_{r}^{\star}}\right)$-$\mathcal{M}_{\text{RIP}}$-strict saddle property. Setting $\epsilon=\frac{1}{160}\sigma_{r}^{\star 3/2}$ completes the proof of the first statement of Proposition 8.